OpenStudy (anonymous):

let f be defined by f(x)= (5x+m, x<2 (-3x^2 + 2m, x>_(greater than or equal to) 2 A? find (in terms of m) lim as x approaches 2+ (f(x))

OpenStudy (loser66):

\(f(x) =\begin{cases}5x+m ~~~~~~x<2\\-3x^2+2m~~~~~~~~~~~~~x\geq 2\end{cases}\)

OpenStudy (loser66):



OpenStudy (anonymous):

\[f(x) = \left\{\begin{array}{rcc} 5x+m &\text{ if} & x <2 \\ -3x^2+2m & \text{if} & x\geq 2 \end{array} \right. \] just seeing if i remembered

OpenStudy (anonymous):

yeah @Loser66 and ok @satellite73

OpenStudy (loser66):

hence, if you need lim when x --> 2+, you need which function? top or bottom?

OpenStudy (loser66):